OpenStudy (anonymous):

The product of three consecutive terms in a geometric sequence is -1000 , and their sum is 15. Find the common ratio. (There are two answers. Separate them by commas) Suggestion: Denote the terms by a/r, a, and ar.

OpenStudy (anonymous):

is this for calculus II?

OpenStudy (anonymous):

this a pretty cool problem, but the hint gives it away

OpenStudy (anonymous):

product is \[\frac{a}{r}\times a\times ar=a^3=-1000\] so you know \[a=-10\]

OpenStudy (anonymous):

then to solve for r you put \[\frac{-10}{r}-10-10r=15\] so \[-\frac{10}{r}-10r=25\] then turn into a quadratic and solve for r

OpenStudy (anonymous):

yeah stupid hint...

OpenStudy (anonymous):

maybe a stupid hint but it sure helped me because i wrote \[a\times ar \times ar^2=-1000\] and got stuck right away, before looking at the hint. reading is a good skill

OpenStudy (anonymous):

I see! So would it be -10, 25? I got -10,5,20 but that isn't it.

OpenStudy (anonymous):

no all we got was \[a=-10\]

OpenStudy (anonymous):

you still have to find "r"

OpenStudy (anonymous):

see my post above as to how to find it

OpenStudy (anonymous):

geometric sequence look like \[(\frac{a}{n})^{z}\] right?

OpenStudy (anonymous):

i got \[r=-\frac{1}{2}\text { or } r=-2\]

OpenStudy (anonymous):

maybe \[\frac{n}{a}\]

OpenStudy (anonymous):

I got r=-2 as well.

OpenStudy (anonymous):

did you solve \[-\frac{10}{r}-10-10r=15\]?

OpenStudy (anonymous):

yep!